If, before any of NaOH, the measured pH is a 2.87 for 0.1 M acetic acid solution calculate the K for acetic acid show calculations.
Answer #1

* of o-1 м Сн соон (44) = 2-87 calculation of Kay ең соон 4) = cңсобсля, вч, Initial no.1 Х Change are + fina) - (o-1-x) ка:pH = 2.87 -2. 87 SHP) = x 10 x = 1. 348 810 M since olor Ž 0.1m -5 ka = at - - 1:8 x 10 0:1